OpenStudy (ashy98):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. A.) y = -x + 6 <--- I think B.) y = x - 6 C.) y = -x - 2 D.) y = x + 2

Nnesha (nnesha):

perpendicular lines have opposite slope for example if your slope is a/b then perpendicular slope suppose to be -b/a flip it and change sign

Nnesha (nnesha):

the answer you choose its same like your question so that mean A iis parallel to the equation

OpenStudy (ashy98):

okay i get its c or d right ?

Nnesha (nnesha):

so plug point into the equation then solve

Nnesha (nnesha):

y = mx+b slope is 1 right so y = x + b now plug point in this equation ( -4 , -2)

Nnesha (nnesha):

did you understand ??

OpenStudy (ashy98):

not really

Nnesha (nnesha):

ok so in order pair you have x and y values like this ( x , y) and your equation is y = x + b now replace x and y by their values then solve for b

Nnesha (nnesha):

can you show your work so i can see where you did mitake ??

OpenStudy (ashy98):

ohhhh its y = x + 2 my bad $_$

Nnesha (nnesha):

yesh :)

Nnesha (nnesha):

because 

Nnesha (nnesha):

to solve for b you have to add 4 both side

Nnesha (nnesha):

then you should get positive 2

OpenStudy (ashy98):

thank you for helping me !!!(:

Nnesha (nnesha):

\(\color{blue}{my :) pleasure }\)